OpenStudy (anonymous):

Jeremy goes to White Water Amusement Park. While there he decides to go down the park's huge water slide called Lightning. If the slide is 48 feet high and the base of the slide is 36 feet from the pool, then what is the length of the slide?

OpenStudy (anonymous):

80^2 + 60^2 = c^2 6400 + 3600 = c^2 10000 = c^2 100 = c So it =100 feet, Hope this helps (✿◠‿◠) .